Not logged in
Fri Nov 30 11:44:57 2007, rob, Configuration, Electronics, ETMX oplev
Fri Nov 30 13:46:50 2007, rob, Configuration, Electronics, ETMX oplev dead again
Fri Nov 30 19:11:05 2007, rana, Configuration, Electronics, ETMX oplev dead again
Fri Nov 30 11:44:57 2007
Reply to this:
In the interests of getting the Xarm alignment script working again, I reset the local damping gains for the test masses to their previous known working values (1), then I noticed that the ETMX oplev was dead. Since the scripts use the oplev motion as a readback for the optic motion, this means the script was basically blindly swinging the optics around. Some monkeying around with swapping HeNe power supplies eventually led to the conclusion that the power strip is funky, since the laser works when plugged into another power strip. Even weirder, the HeNe and the power supply indicator light have some sort of XOR relationship going on. When one works, the other doesn't. Steve will sort out this confusion later; we're good for now.